OpenStudy (anonymous):

WILL FAN AND MEDAL PLEASE HELP The wavelength of red light is given as 6.5 × 10È¥ m. If another wave of an unknown type has a frequency of 5.3 × 10¹¢ Hz. What is the relation between the wavelength of the unknown wave and red light? (note: c = 2.998 × 10§ m/s) The wavelength of the unknown wave is less than the wavelength of the red light the wavelength of the wave is greater than the wavelength of the red light the wavelength of the wave is equal to the wavelength red light the wavelength of the wave is not related to the wavelength of the red light

OpenStudy (rushwr):

\[c = f \lambda \]

OpenStudy (rushwr):

c = velocity f= frequency \[\lambda = wavelength \]

OpenStudy (anonymous):

I have no clue what to do with that...

OpenStudy (anonymous):

I think it's A but I'm not sure...

OpenStudy (rushwr):

to compare the wavelengths we have to obtain a value for the wavelength of the unknown wave using the above equation obtain a value. \[\lambda = \frac{ c }{ f } \]

OpenStudy (anonymous):

So... C? because that's what's looking right now...

OpenStudy (rushwr):

no no !! Wait I'll tell you but tell me the value for C !!! It's a little unclear !

OpenStudy (rushwr):

if u don't mind right the values here again it's a little unclear !! type it here, don't copy paste

OpenStudy (anonymous):

2.998*10^8 m/s

OpenStudy (anonymous):

sorry open study turns the notations into weird symbols

OpenStudy (rushwr):

what about the values for others !!! It's not clear ! ^_^

OpenStudy (anonymous):

6.5*10^-7 is the first one and 5.3*10^15

OpenStudy (rushwr):

aaah okai. Now we know the wavelength of the red light. We are given with the frequency of the unknown wave. Looking at the answers we can see that they have compared the wavelengths. We only know the wavelength of the red light, to compare the wavelengths we have to find the wavelength of the unknown wave right? Do u understand up to this point ?

OpenStudy (anonymous):

I do.

OpenStudy (anonymous):

I just don't know how to find the wave length, I get confused with the numbers

OpenStudy (rushwr):

Now we have to find the wavelength of the unknown wave. So we are gonna use the above equation: \[c = f \lambda \] \[\lambda = \frac{ c }{ f } = \frac{ 2.998 * 10^{8} }{ 5.3 * 10^{15} }\]

OpenStudy (rushwr):

\[\lambda = 5.65 * 10^{-8}\]

OpenStudy (rushwr):

As we can see that the unknown wave has a shorter wavelength than the red light !

OpenStudy (anonymous):

Thank you!